OpenStudy (anonymous):

,For the given functions f and g, find the requested composite function value. f(x) = /sqrt{x + 5} , g(x) = 2x ; Find (f ∘ g)(2). Seleccione una respuesta. a. 2 {7} b. 3 c. {14} d. 2{14}

OpenStudy (anonymous):

is it D the right answer?

OpenStudy (anonymous):

fog is sqrt(2x+5)

OpenStudy (anonymous):

Who knows the right answer?

OpenStudy (anonymous):

the answer is 3 ,

OpenStudy (anonymous):

\[f \circ g(x) = \sqrt{2x+5}\] \[f \circ g (2) = \sqrt{2 \times 2 +5}=\sqrt9=3\]

OpenStudy (anonymous):

C

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B